OpenStudy (anonymous):

What is 220% of 400?

OpenStudy (bahrom7893):

400 - 100% x - 220% x = (220*400) / 100 x = 220*4 = 880. Now try something on you own

OpenStudy (bahrom7893):

find 210% of 400

OpenStudy (anonymous):

hahahaha yeah i get the wrong answer all the time when i think i got the right answer

OpenStudy (anonymous):

i just tried 5 problems and im still getting them wrong

OpenStudy (anonymous):

im trying to folllow your steps in hopes to actually get 1 right at least

OpenStudy (anonymous):

isnt it 840